# Pareto-Optimal Learning from Preferences with Hidden Context

### v1

@@ 5.2 Lexicase Selection
To obtain a set of Pareto optimal policies, we adopt the idea of lexicase selection [51, 32], which performs selection based on a lexicographic ordering of metrics. These metrics are shuffled into a random order (and therefore a random prioritization) for each selection event, and only the candidates that perform exactly the best on each metric are kept around. This process is repeated until we run out of metrics or are left with a single candidate individual.
@@
A key property of lexicase selection is that it selects candidates that are Pareto-optimal relative to a starting set of candidates that tend to spread the corners of the Pareto front and thus be diverse [39]. This idea has been utilized in many machine learning optimization problems for improving generalization, as shown in recent work [21, 22, 44, 8, 20]. In the context of preference-based reward learning, preferences are used as binary metrics that can be generated in one of two ways: either a pass/fail approach or a likelihood approach.
In the pass/fail approach, each metric is passed if the candidate policy correctly ranks the pair of segments in the preference corresponding to that metric. Note that the content of the preference does not change the chance of passing a metric, and so this framework does not take into account Boltzmann rationality, where the probability of giving a preference depends on the difference in the exponentiated utility between the two alternatives. To account for the fact that people might give noisy preferences within their hidden context group, we use a likelihood passing approach, where a policy passes a metric with probability determined by the likelihood of that preference under the policy, with a specific regularization mechanism.
An important quality presented by lexicase selection is the notion that selected individuals are assessed with respect to their current pool. If all candidates in the pool fail a given preference (either deterministically or in the stochastic selection approach), then no individuals are filtered out. This idea helps overcome the fact that there might be contradictory preferences found in the dataset. When this happens, preferences that come earlier in the shuffle take priority over those that come later. For example, if the preference that comes first in the shuffle is contradictory to the preference that comes second, no individuals that pass the first preference will pass the second, and so no individuals will be filtered out (equivalent to the second preference not being in the shuffle). With the random orderings that lexicase selection uses, different preferences will be prioritized at different times, selecting policies or reward functions that pass diverse sets of compatible preferences.
It is important to note that lexicase selection does not guarantee that the individuals are globally pareto-optimal (i.e. pareto-optimal with respect to every policy or reward function possible). Instead, it guarantees that the individuals are Pareto-optimal with respect to the other candidates in the same pool. Whilst this distinction is important, we do not believe it significantly degrades our ability to recover the optimal policy for each hidden context group. This is because, with a large enough population, lexicase selection is able to continually search for individuals on the Pareto-frontier through successive generations.
@@ 6 Experiments
In this section, we detail our experimental results to validate the proposed POPL method. Further implementation details are provided in Appendix C.
6.1 Baselines
Throughout our experiments, we will use 3 main baselines. In the experiments on reward function inference, we use Bayesian Reward Extrapolation (B-REx) [12] as a baseline, as it generates a large set of reward function hypotheses based on a Boltzmann-rational likelihood function, and has demonstrated efficacy in RL domains.
For our policy inference experiments, we compare to Contrastive Preference Learning [31] as it is a leading RLHF algorithm for arbitrary sequential tasks. We also use a naive method of learning a set of policies based on CPL that we call Multi-CPL. In this approach, after pretraining, we fine-tune the last layer using the CPL objective multiple times to generate a large set of reward functions. Although we could do full network fine-tuning, we wanted to hold constant the trainable parameters available to each approach to ensure a fair comparison to POPL, which only uses last-layer fine-tuning. A secondary reason for the policy inference experiments is to verify whether our proposed technique is sensitive to the choice of preference generation model (regret vs. partial return).
For our language model (contextual bandit) experiments, we compare to both B-REx and Distributional Preference Learning (DPL) [50], as well as standard the standard RLHF paradigm [17], as these present a variety of approaches for generating reward models that can be used to ensure fairness across groups.
6.2 Metrics
Given a set of reward functions or policies, we can verify how well they perform on the two downstream tasks we have identified for this work: personalization and fairness. For personalization, we inspect the content of the personalized policies or reward functions to verify their alignment with each hidden context group’s preferences. For fairness, we measure how well a new policy $\pi_{\text{fair}}$ performs with respect to all hidden context groups. Given a set of policies $\Pi$, one can generate $\pi_{\text{fair}}$ by picking actions that have the highest lower (10%) quantile probability across the set. In other words, we ensure that no single group is having its values undermined (by taking low-probability actions) in an attempt to satisfy a different group. Although this is a relatively simple notion of fairness, this method could be extended to be compatible with other fairness optimization approaches [42].
@@ 6.3 Synthetic Stateless Experiment
(a) B-REx Distribution
(b) POPL Final Set
(c) B-REx Catering
(d) POPL Catering
Figure 3: (a) and (b) shows all the reward functions that are the output of the reward inference methods. (c) and (d) show the catered reward functions for each of the two hidden context groups $z=0$, $z=1$. To ensure the reward functions are comparable, for all these figures, the $y$-axis shows how the underlying reward functions rank the 100 differing state interpolations ($x$-axis).This is monotonically equivalent to the utility assigned to each state by the reward functions. From a set of reward functions that is inferred from a diversity of human preferences, we select a single reward function for each unique group with a small number of preferences (2% the size of the training set). POPL is able to cater for both groups, while B-REx is only able to cater for one of the two groups ($z=1$, red line). For B-REx, the $z=0$ (green) group’s catered reward function doesn’t capture the fact that any state $a<0.8$ is preferred to any state $a\geq 0.8$.
Following the synthetic experiments outlined by Siththaranjan et al. [50], we compare B-REx and lexicographic reward inference methods on learning a distribution of reward functions based on preferences from two distinct groups of humans defined by a hidden context variable $z\sim\mathcal{B}(0.5)$ where $\mathcal{B}(0.5)$ is a Bernoulli distribution. We compare the distributions based on their representation of the various hidden context groups that might have different preferences.
The utility in this scenario can be modeled as
@@
Figure 3 shows the resulting reward functions from performing reward inference on a dataset of preferences with underlying hidden context using POPL and B-REx. For each of these methods, we display all the reward functions created through each approach. As there are two equally proportioned groups, each with overlapping preferences, B-REx tends to concentrate all of the probability mass on a single one of these groups. POPL, on the other hand, is able to learn two distinct equivalence classes of reward functions. The reward functions inferred are monotonically increasing in the range $a\in[0,0.8]$. After this range, where the hidden context has an effect, some of the reward functions predict a decreased utility to reflect the first group’s preferences, and other reward functions predict increased utility to predict the other group’s preferences.
In order to test whether POPL achieves our goal of having at least one reward function in the set that covers each hidden context group, we inspect some selected reward functions for each group. We use a smaller set of the preferences that all have a shared hidden context, and attempt to select a reward function from the distribution that is catered to a user in this group. Figure 3 also shows the results of catering a reward function for each of the hidden context classes $z=0$ and $z=1$. Due to B-REx using the Boltzmann rationality assumption, it concentrates much of the distribution on the $z=1$ case, and does not capture the preferences given by the $z=0$ group. POPL, on the other hand, is able to recover the reward functions for both groups from the learned distributions.
@@ 6.4 Minigrid Policy Inference
Now that we have demonstrated POPL’s efficacy in reward inference from preferences with hidden context, we perform a second set of experiments to verify whether 1) POPL is able to generate policies directly from preferences, and 2) POPL is able to perform in a sequential RL domain, where annotators’ hidden context is persistent (i.e. potentially affects more than one segment preference annotation).The domain used in these experiments is outlined in Figure 4(a). The hidden context groups in this scenario delineate whether the annotator inherently prefers the bottom or top door to be used to get to the goal. The preferences were labeled according to the regret preference model from members of both groups (extracted from the optimal policy for each group’s ground truth reward model). After running this optimization, catered policies for each group can be found in Figures 4(c) and 4(d) for POPL, and 4(e) and 4(f) for MultiCPL. We find that POPL is able to successfully cater policies for both groups of people, despite not having labels regarding their group membership.
@@
(b) POPL for Group 1
(c) POPL for Group 2
(d) Hidden Context Groups
@@
Figure 4: (a) An example domain where hidden context is present. The agent (red triangle) must make it to the solid green goal tile as fast as possible. The agent must choose one of the two doors (top or bottom) to use to reach the goal. (d) Depending on their identity, a human might prefer a policy that uses the top door over the bottom door, or vice versa. (b) and (c) show the aggregate state occupancy of the policies personalized for each group using POPL. (e) and (f) show the state occupancy of the policies personalized for each group using MultiCPL. Although MultiCPL has the same number of policies generated as POPL, the set that is outputted does not contain the optimal policy for Group 1. The set of policies generated by POPL contains the optimal policy for both groups, without having access to annotator identity labels.
@@ 6.5 Language Model Experiments
A final set of experiments that we perform tests the ability of POPL to scale to domains involving human annotations. We investigate whether POPL can be sensitive to hidden context in whether annotators prefer harmless or helpful responses [1]. When reward models are trained on the entire set of preferences, whether they were generated based on helpfulness or harmlessness is hidden context, as this information affects preferences but is not available to a reward inference system.
Importantly, preferences based on helpfulness and harmlessness can often be contradictory. In fact, Wei et al. [53] find examples of user prompts that directly pit these objectives against each other, leading a language model to output harmful outputs, a phenomenon known as jailbreaking. An RLHF system built with hidden context in mind would help detect jailbreaking before a harmful output would be given to a user. In the context of this work, a model that is susceptible to jailbreaking would be unfair to certain groups (compromising its efficacy for the harmlessness group in order to optimize for the helpfulness group).
Table 1 presents the jailbreak rates and helpfulness accuracy for standard RLHF, B-REx, DPL, and our proposed POPL. For B-REx and POPL, we generate a set of reward functions by extrapolating the last layer of a fine-tuned Llama-2-7b [52] preference model. Default settings use the mean reward across the entire set. For fairness optimization, we use ${}^{\text{th}}$ percentile of reward values across all the reward functions in the set.
@@ 7 Conclusion
When learning from human preferences for the sake of aligning to human values, one must first decide which humans to pay attention to. As these systems rely on point estimates of return or regret, they can at most align with a single human’s preferences. We have formalized this as the problem of reinforcement learning from human preferences with hidden context. Under this conception, a set of policies must be generated that contains the optimal policy for each group of people.
@@
We verified that lexicase selection can be used to generate diverse distributions of either reward functions or policies that align with the diverse preferences that human annotators have. We evaluated and verified the performance of POPL in a variety of domains, including a synthetic stateless domain, a sequential RL domain, and even language model jailbreak detection. Across these domains, we have demonstrated POPL’s efficacy when compared to contemporary algorithms in dealing with hidden context in the preferences.
One major limitation of this work is the lack of use of gradients in training policies. The optimization procedure used after lexicase selection relies on random variations and repeated selections, which allows for effective trade-offs between exploration and exploitation of the preference landscape. However, it may be possible to augment the core idea in future work to allow it to utilize gradients. Furthermore, a study into the conditions required for the output of the procedure to be globally Pareto-optimal could be instrumental.
Acknowledgments
This work has taken place in the Safe, Correct, and Aligned Learning and Robotics Lab (SCALAR) at The University of Massachusetts Amherst. SCALAR research is supported in part by the NSF (IIS-2323384), AFOSR (FA9550-20-1-0077), and the Center for AI Safety (CAIS).
LD and LS were supported by the National Science Foundation under Grant No. 2117377. This work was performed in part using high-performance computing equipment at Amherst College obtained under National Science Foundation Grant No. 2117377. Any opinions, findings, and conclusions or recommendations expressed in this publication are those of the authors and do not necessarily reflect the views of the funding agencies. The authors would like to thank Philip Thomas, members of the SCALAR lab at UMass Amherst, members of the PUSH lab at Amherst College, University of Massachusetts Amherst, and Hampshire College for helpful discussions, and anonymous reviewers for their thoughtful feedback.
Author Contributions
RB conducted preliminary experimentation, conceived and framed the core ideas, implemented the core algorithm, ran the synthetic and Minigrid experiments, and generated the figures. RB also contributed significantly to the writing and organization of the manuscript. LD implemented the algorithm in an efficient framework optimized for scalability, conducted the LLM experiments, and formulated the associated analyses. LD also contributed to the ideation and provided feedback on multiple drafts of the manuscript.
LS provided oversight on the high-level conceptualization of lexicase selection and its application. LS offered continuous advice, reviewed the manuscript, and contributed to the theoretical framing of the work. SN oversaw the project from inception to completion, contributed significantly to ideation and conceptual framing, and ensured the project stayed aligned with its broader goals. SN also helped with securing funding and provided extensive feedback on the manuscript.
@@ Appendix A Full Algorithm
Algorithm 1 One Step of Pareto Optimal Preference Learning
@@ Appendix B Proofs
Proof of Theorem 1 (Contradiction). Let us assume that there is a policy $\pi^{*}_{z}$ that is the optimal policy according to a hidden context group $z$. This means $\pi_{z}$ passes all the preferences compatible with the values of group $z$ and fails only the preferences that are not compatible with preferences given by group $z$. For sake of contradiction, we assert that this reward function is not Pareto optimal with respect to all other reward function candidates. This means there exists another reward function $\pi^{\prime}$ that performs better than or equal to $\pi^{*}_{z}$ across every preference in $\mathcal{P}$, including those generated by group $z$. This is a clear contradiction as that would imply $\pi^{\prime}$ is the optimal policy for group $z$ instead of $\pi_{z}$.

### v2

@@ 5 Pareto Optimal Preference Learning
To obtain a set of Pareto optimal policies, we adopt the idea of lexicase selection (Spector, 2012; Helmuth et al., 2015), which uses a random ordering of metrics for each selection event, with only the candidates that perform the best on each successive metric retained for filtering by the remaining metrics. This process is repeated until all metrics are exhausted or a single individual remains. Through this process lexicase selection prioritizes, over multiple selection events, each particular metric and each particular combination of metrics to the exclusion of all others. We can consider this to be a “particularity” approach for achieving pluralistic outcomes (Spector et al., 2024).
In the setting of preference learning, each metric corresponds to a preference sourced from a human with hidden context. The preference is ‘passed’ if the candidate policy correctly ranks the pair of segments in the preference corresponding to that metric (formally defined in Section 4). If no individuals in the current pool pass the preference, all individuals make it through this selection step. With this feature, contradictory preferences are addressed by giving priority to the first preference in the shuffle. Each random shuffle of preferences therefore results in a diverse profile of reward functions being selected for. Figure 3 shows an example of a single selection event.
@@
A key property of lexicase selection is that it selects candidates that are Pareto-optimal relative to a starting set of candidates (as opposed to all possible candidates). These individuals tend to spread the corners of the Pareto front and thus be diverse (La Cava et al., 2019). Lexicase selection also gives individuals that are good at more subsets of things greater weight. This idea has been utilized in many machine learning optimization problems for improving generalization, as shown in recent work (Ding et al., 2022; 2023; Ni et al., 2024; Boldi et al., 2023; Ding & Spector, 2022).
Overview
With a method to select Pareto-optimal candidates such as lexicase selection in hand, one can infer a set of reward functions or policies directly from preferences. Initially, a random set of candidate models is created. Then, the chosen method is applied to select (with replacement) the Pareto-optimal candidates from this random starting set. This pool is perturbed by adding random Gaussian noise, generating a new set of candidates. The selection and perturbation steps are repeated iteratively until the average performance converges, or a fixed number of iterations is passed. The final set of candidates should align with the preferences of hidden context groups. A full overview of our algorithm can be found in Appendix A.
@@ 6 Experiments
In this section, we detail our experimental results to validate the proposed POPL method. To verify that POPL can work in a large variety of settings at different scales, as well as for generating both reward functions and policies, we perform four sets of experiments. A synthetic, stateless experiment (reward inference), a Minigrid RL environment (policy inference), a Metaworld robotics environment (policy inference) and LLM finetuning from human preferences (reward inference). Further implementation details are provided in Appendix C.
Baselines
Throughout our experiments, we will use 3 main baselines. In the experiments on reward function inference, we use Bayesian Reward Extrapolation (B-REx) (Brown et al., 2020b) as a baseline, as it generates a large set of reward function hypotheses (i.e., candidate models) based on a Boltzmann-rational likelihood function, and has demonstrated efficacy in RL domains. For our policy inference experiments, we compare to Contrastive Preference Learning (Hejna et al., 2024) as it is a leading RLHF algorithm for sequential tasks. We also use a naive method of learning a set of policies based on CPL that we call Multi-CPL. In this approach, after pretraining, we fine-tune the last layer using the CPL objective multiple times to generate a large set of policies. Although we could do full network fine-tuning, we wanted to hold constant the trainable parameters available to each approach to ensure a fair comparison to POPL, which uses last-layer fine-tuning. Policy learning settings in this work model human preferences as being generated based on regret, as opposed to partial return (Knox et al., 2022). Including the policy inference experiments allows us to ensure our method is not sensitive to assumptions regarding how the preferences are generated. For our language model (contextual bandit) experiments, we compare to both B-REx and Distributional Preference Learning (DPL) (Siththaranjan et al., 2023), as well as standard the standard RLHF paradigm (Christiano et al., 2017), as these present a variety of approaches for generating reward models that can be used to ensure fairness across groups.
Metrics
Given a set of reward functions or policies, we can verify how well they perform on the two downstream tasks we have identified for this work: personalization and fairness. For personalization, we inspect the content of the personalized policies or reward functions to verify their alignment with each hidden context group’s preferences. For fairness, we ensure that no single group is having its values undermined (by taking low-probability actions) in an attempt to satisfy a different group. Although this is a relatively simple notion of fairness, this method could be extended to be compatible with other fairness optimization approaches Mehrabi et al. (2021).
@@ 6.1 Synthetic Stateless Experiment
(a) B-REx Catering
(b) POPL Catering
Figure 4: (a) and (b) show the catered reward functions for each of the two hidden context groups $z=0$, $z=1$. From a set of reward functions that is inferred from a diversity of human preferences, we select a single reward function for each unique group with a small number of preferences (2% the size of the training set). POPL is able to cater for both groups, while B-REx is only able to cater for one of the two groups ($z=1$, red line). For B-REx, the $z=0$ (green) group’s catered reward function doesn’t capture the fact that any state $a<0.8$ is preferred to any state $a\geq 0.8$.
@@
Following the synthetic experiments outlined by Siththaranjan et al. (2023), we compare B-REx and POPL on learning from preferences where with hidden context variable $z\sim\mathcal{B}(0.5)$ where $\mathcal{B}(0.5)$ is a Bernoulli distribution. The utility in this scenario can be modeled as
@@
In order to test whether POPL covers the hidden context groups, we inspect some selected reward functions for each group. We use a smaller set of the preferences that all have a shared hidden context, and select a reward function for each group. Figure 4 shows the results of catering a reward function for each of the hidden context classes $z=0$ and $z=1$. Due to B-REx using the Boltzmann rationality assumption, it concentrates much of the distribution on the $z=1$ case, and does not capture the preferences given by the $z=0$ group. POPL, on the other hand, is able to recover the reward functions for both groups from the learned distributions.
@@ 6.2 Minigrid Policy Inference
(b) HC Groups
(c) POPL for Group 1
(d) POPL for Group 2
@@
Figure 5: Minigrid experiments. Plots in (c), (d), (e) and (f) show average state occupancy for policies catered for each hidden context group. POPL is able to cater distinct policies for each group, while MultiCPL collpases to a single group’s preferences.
After demonstrating POPL’s efficacy in reward inference from preferences with hidden context, we perform a second set of experiments to verify whether 1) POPL is able to generate policies directly from preferences, and 2) POPL is able to perform in a sequential RL domain, where annotators’ hidden context is persistent (i.e. potentially affects more than one segment preference annotation). The domain used in these experiments is outlined in Figure 5(a). The agent (red triangle) must make it to the solid green goal tile as fast as possible. The agent must choose one of the two doors (top or bottom) to use to reach the goal. The hidden context groups in this scenario delineate whether the annotator inherently prefers the bottom or top door to be used to get to the goal (Figure 5(b)). The preferences were labeled according to the regret preference model from members of both groups (extracted from the optimal policy for each group’s ground truth reward model).
After running this optimization, the state occupancy distribution for catered policies for each group can be found in Figures 5(c) and 5(d) for POPL, and 5(e) and 5(f) for MultiCPL. We find that POPL is able to successfully cater policies for both groups of people (as exhibited by policies reaching the goal via both doors), despite not having labels regarding their group membership. MultiCPL, on the other hand, is unable to cater a policy for Group 1.
(a) button-press-v2
(b) reach-v2
Figure 6: Metaworld Policy Inference Results. Box plots outline the performance of the best (catered) individual from each population on both identities across 10 random seeds. We also show the average performance of the best “compromise,” or the single policy that does the best across both identities. POPL tends to have a higher catered policy performance across both identities, and also discovers a more fair compromise between values of the two groups.
6.3 Metaworld Policy Inference
In order to verify how well POPL can infer policies in larger scale sequential environments, we include results performing policy inference on the Metaworld Robotics Benchmark (Yu et al., 2019). We artificially create two hidden context groups, one that prefers safe (low angular velocity) robotic movements, and one that prefers speed (having tasks completed quickly). We generate preferences from these two groups at random, and then compare POPL and MultiCPL’s ability to cater individual policies for each group.
Figure 6 outlines the performance of all the policies generated by POPL and the MultiCPL baseline. We also include a case study comparing a single run of POPL and MultiCPL in Figure 7. POPL is able to generate policies that outperforms the MultiCPL and behavior cloning baselines. POPL finds policies that are maximally good for either group, as well as those that find strong compromises between the two group’s values.
Figure 7: Case study for a single button-press-v2 run. POPL finds policies that perform well under both ground truth reward functions. We also include a behavior cloning (BC) baseline, where the policies are simply trained to match the demonstrations.
@@ 6.4 Language Model Experiments
In this section, we test the ability of POPL to scale to domains involving human annotations. We investigate whether POPL can be sensitive to hidden context in whether annotators prefer harmless or helpful responses (Bai et al., 2022). When reward models are trained on the entire set of preferences, whether they were generated based on helpfulness or harmlessness is hidden context, as this information affects preferences but is unavailable to a reward inference system.
Importantly, preferences based on helpfulness and harmlessness can often be contradictory. In fact, Wei et al. (2024) find examples of user prompts that directly pit these objectives against each other, leading a language model to output harmful outputs, a phenomenon known as jailbreaking. An RLHF system built with hidden context in mind would help detect jailbreaking before a harmful output would be given to a user. In the context of this work, a model that is susceptible to jailbreaking would be unfair to certain groups (compromising its efficacy for the harmlessness group in order to optimize for the helpfulness group).
Table 1 presents the jailbreak rates and helpfulness accuracy for standard RLHF, B-REx, DPL, and our proposed POPL. For B-REx and POPL, we generate a set of reward functions by extrapolating the last layer of a fine-tuned Llama-2-7b (Touvron et al., 2023) preference model. Default settings use the mean reward across the entire set. For fairness optimization, we use ${}^{\text{th}}$ percentile of reward values across all the reward functions in the set.
@@ 7 Conclusion
When learning from human preferences for the sake of aligning to human values, systems often rely on point estimates of return or regret, limiting them to aligning to a single group of humans. Preferences, however, often come from distinct groups with diverse preferences. We have formalized this as the problem of preference learning with hidden context. Under this conception, a set of policies must be generated that contains the optimal policy for each distinct group.
@@
We verified that lexicase selection can be used to generate diverse distributions of either reward functions or policies that align with the diverse preferences that human annotators have. We evaluated and verified the performance of POPL in a variety of domains, including a synthetic stateless domain, a Minigrid RL domain, a Metaworld Robotics benchmark, and even language model jailbreak detection. Across these domains, we have demonstrated POPL’s efficacy when compared to contemporary algorithms in dealing with hidden context in the preferences. Without modifications to the framework, POPL can be used to optimize for diverse reward functions or policies, and can work in stateless and sequential domains at a variety of scales.
One limitation of this work is the lack of use of gradients in training policies. The optimization procedure used after lexicase selection relies on random variations and repeated selections, which allows for effective trade-offs between exploration and exploitation of the preference landscape. Although empirically verified to work well, it may be possible to augment the core idea in future work to allow it to utilize gradients. Furthermore, a study into the conditions required for the output of the procedure to be globally Pareto-optimal could be instrumental.
Reproducibility Statement
We are committed to the reproducibility of our results. We include full code to reproduce the results in this paper as supplemental material. This code includes dataset generation and the full POPL training pipeline. Furthermore, we outline experimental details needed to independently reproduce the results in Appendix C. The theory performed in Section 4 has proofs associated in Appendix B and assumptions outlined therein.
@@ Appendix A Full Algorithm
Algorithm 1 Pareto Optimal Preference Learning
@@ Appendix B Proofs
Proof of Theorem 1 (Contradiction).
Let us assume that there is a policy $\pi^{*}_{z}$ that is the optimal policy according to a hidden context group $z$. This means $\pi^{*}_{z}$ passes all the preferences compatible with the values of group $z$ and fails only the preferences that are not compatible with preferences given by group $z$. For sake of contradiction, we assert that this reward function is not Pareto optimal with respect to all other reward function candidates. This means there exists another reward function $\pi^{\prime}$ that performs better than or equal to $\pi^{*}_{z}$ across every preference in $\mathcal{P}$, including those generated by group $z$. This is a clear contradiction as that would imply $\pi^{\prime}$ is the optimal policy for group $z$ instead of $\pi^{*}_{z}$.